# Attribute-driven Disentangled Representation Learning for Multimodal Recommendation

### v1

@@ 3.2.1. High-Level Attribute-driven Disentangled Representation Learning
Intra-Modality Disentanglement. After extracting the feature vector from each modality, we simply mechanically split the vector into several chunks. Nevertheless, different attribute factors are still entangled within the chunks. That is, for a chunk, it may contain both price- and brand-related features. Therefore, in order to disentangle attribute factors in each modality feature, we employ an attribute classifier to force each chunk to be sufficient to predict the corresponding attribute, as shown in Fig. 1(a). Taking the $k$-th chunk $\bm{q_{t}^{k}}$ of item textual embedding as an example,
(3) $\begin{cases}\mathbf{z}^{k}&=\mathbf{W}_{intra,t}\bm{q_{t}^{k}}+\mathbf{b}_{intra,t},\\ l_{intra,t}^{k}&=-\sum_{n=1}^{K}\tilde{z}_{n}^{k}\log\frac{\exp{z_{n}^{k}}}{\sum\nolimits_{m}\exp{z_{m}^{k}}}\end{cases}$
where $\mathbf{W}_{intra,t}\in\mathbb{R}^{K\times\frac{d}{K}}$ and $\mathbf{b}_{intra,t}\in\mathbb{R}^{K}$ are the weight matrix and bias vector of the classifier, respectively. $\mathbf{z}^{k}$ is the predicted logits and $z_{n}^{k}\in\mathbf{z}^{k}$. $\tilde{z}_{n}^{k}$ denotes the ground truth (attribute label) of chunk $\bm{q_{t}^{k}}$. For all the chunks in textual embedding $\bm{q_{t}}$, we have the following loss for all attribute factors in textual features:
@@
Inter-Modality Disentanglement. In addition to separately disentangling each modality, a serious challenge in disentangling representation learning for multi-modal features is handling the inter-relationship between the factors disentangled from multiple modalities. Intuitively, the chunks of the same attribute factor from different modality features should be consistent. For example, for the brand attribute, an item should have the same brand information in both the visual and textual features.
Inspired by this, we disentangle attribute factors in different modality features from the perspectives of consistency in multi-modality recommendations. Two chunks of the same attribute factor from two different modality features should have a high similarity because they are related to the same attribute. In contrast, two chunks from different modality features of different attribute factors should have low similarity as they provide different attribute information.
To this end, we design a cross-modal contrastive loss as shown in Fig. 1(b). Specifically, for any two modality representations, such as $\bm{q_{t}}$ and $\bm{q_{v}}$, we take two chunks of the same attribute factor (i.e., $(\bm{q_{t}^{k}},\bm{q_{v}^{k}})$) as positive pairs, and two chunks corresponding to different attributes (i.e., $(\bm{q_{t}^{k}},\bm{q_{v}^{n\neq k}})$, $(\bm{q_{t}^{n\neq k}},\bm{q_{v}^{k}})$) as negative pairs. After that, the cross-modal contrastive loss between textual and visual features is defined as follows:
(6) $\begin{cases}l^{k}_{t\rightarrow v}&=-\log\frac{\exp((\bm{q_{t}^{k}}\cdot\bm{q_{v}^{k}})/\tau)}{\sum_{n=1}^{K}\exp((\bm{q_{t}^{k}}\cdot\bm{q_{v}^{n}})/\tau)},\\ l^{k}_{v\rightarrow t}&=-\log\frac{\exp((\bm{q_{v}}^{k}\cdot\bm{q_{t}^{k}})/\tau)}{\sum_{n=1}^{K}\exp((\bm{q_{v}^{k}}\cdot\bm{q_{t}^{n}})/\tau)},\\ l_{t\leftrightarrow v}&=\sum_{k=1}^{K}l^{k}_{t\rightarrow v}+\sum_{k=1}^{K}l^{k}_{v\rightarrow t},\end{cases}$
@@
3.2.2. Low-level Attribute-driven Disentangled Representation Learning.
Since each factor has a clear attribute meaning, each attribute should have its corresponding specific attribute values. Specifically, each attribute $k$ is associated with a list of possible attribute values $(\tilde{y}^{k}_{1},\tilde{y}^{k}_{2},\cdots,\tilde{y}^{k}_{A_{k}})$, where $A_{K}$ is the total number of possible values for that attribute. For example, the attribute value of popularity can be stratified into five distinct levels: Super Popular, Popular, Moderate, Emerging, and Unknown. We hope that the representations disentangled based on attributes can reflect the specific attribute values of the item.
It is obvious that any single modality may not contain sufficient information for a particular attribute. For example, images may more intuitively reflect attributes such as color and brand of an item, but may not directly reflect attributes such as price or popularity. Therefore, in order to comprehensively and accurately depict the attribute values of the item, we first integrate the features from all modalities together. To achieve this, we apply a multimodal attention mechanism to measure the emphasis of different modalities on different attributes. Specifically, for the $k$-th attribute, the specific attention values to different modalities are computed by a two-layer neural network:
(8) $\begin{cases}\hat{\mathbf{a}}^{k}=\textbf{W}_{a2}\tanh(\mathbf{W}_{a1}(\bm{q_{i}^{k}}+\bm{q_{t}^{k}}+\bm{q_{v}^{k}})+\bm{b}_{a}^{k}),\\ \mathbf{a}^{k}=Softmax(\hat{\mathbf{a}}^{k}),\end{cases}$
where $\mathbf{W}_{a1}\in\mathbb{R}^{3\times\frac{d}{K}}$ and $\textbf{W}_{a2}\in\mathbb{R}^{3\times 3}$ are the weight matrices corresponding to the first and second layers of the neural network, respectively. $\bm{b}_{a}^{k}$ denotes the bias vector and $\tanh$ is the activation function. $Softmax$ is adopted to normalize $\hat{\mathbf{a}}^{k}$ to a probability distribution.
@@
(9) $\bm{q_{y}^{k}}=a_{i}^{k}\cdot\bm{q_{i}^{k}}+a_{t}^{k}\cdot\bm{q_{t}^{k}}+a_{v}^{k}\cdot\bm{q_{v}^{k}},$
where $a_{i}^{k}$, $a_{t}^{k}$, $a_{v}^{k}\in\mathbf{a}^{k}$ are the attention weights of item ID embedding, textual feature and visual feature, respectively.
Similar to the disentangling at the intra-modality disentanglement in Section 3.2.1, we aim to achieve disentangled representation learning at the low level through attribute value prediction via a classification layer (as shown in Fig. 1(c)):
(10) $\begin{cases}\mathbf{y}^{k}&=\mathbf{W}_{k}\bm{q_{y}^{k}}+\mathbf{b}_{k},\\ l^{k}_{low}&=-\sum_{i=1}^{A_{K}}\tilde{y}_{i}^{k}\log\frac{\exp{y_{i}^{k}}}{\sum\nolimits_{j}\exp{y_{j}^{k}}},\\ l_{low}&=\sum_{k=1}^{K}l^{k}_{low}\end{cases}$
where $\mathbf{W_{k}}\in\mathbb{R}^{K\times\frac{d}{K}}$ and $\mathbf{b}_{k}\in\mathbb{R}^{K}$ are the weight matrices and bias vector of the classifier, respectively. We supervise the training of each attribute subspace via independent attribute classification tasks defined in the form of a cross-entropy loss.
@@ 3.3.1. Preference Prediction
So far we have discussed how to obtain attribute-driven disentangled representations $\bm{p_{u}}=(\bm{p_{u}}^{1},\bm{p_{u}}^{2},\cdots,\bm{p_{u}}^{K})$ and $\bm{q_{y}}=(\bm{q_{y}^{1}},\bm{q_{y}^{2}},\cdots,\bm{q_{y}^{K}})$ for the user $u$ and item $i$, respectively. By assigning each disentangled representation with an attribute as described above, we are able to predict users’ preferences at the attribute level. This enhances the interpretability and controllability of our model. Specifically, to estimate a user $u$’s preference for an item $i$, it is crucial to consider her preference for each attribute of the item. To achieve this, we first compute the user’s preference score for each attribute of the item, and subsequently, integrate the scores of each attribute to estimate the user’s overall preference for the item:
(11) $\begin{cases}s_{u,i,k}=\sigma(\bm{p_{u}^{k}}\cdot\bm{q_{y}^{k}}),\\ s_{u,i}=\sum_{k=1}^{K}s_{u,i,k},\end{cases}$
@@ 3.3.2. Training Protocol
Based on the predicted preference score above, we recommend a set of top-$n$ ranked items which match the target user’s preferences. The Bayesian Personalized Ranking (BPR) loss function (Rendle et al., 2009) is employed to optimize the model parameters $\bm{\Theta}$,
@@
where $\lambda$ is the coefficient controlling $L_{2}$ regularization; $\mathcal{D}$ denotes the training set; $i_{+}$ and $i_{-}$ are the observed and unobserved item in the interaction records of user $u$, respectively. Overall, the total loss of AD-DRL is formulated as,
@@
where $\alpha$, $\beta$ and $\gamma$ are the hyperparameters to control the weight of three disentanglement module.
@@ 4. Experiments
We conducted extensive experiments on three benchmark datasets to validate the effectiveness of the proposed method. In particular, the experiments are mainly performed to address the following research questions:
  • RQ1: Can the proposed disentangled representation learning method outperform existing state-of-the-art models?
  • RQ2: Can AD-DRL effectively learn disentangled representations by utilizing attributes?
  • RQ3: How do the controllability and interpretability of our proposed AD-DRL perform?
  • RQ4: How do the key components of AD-DRL, specifically high- and low-level attribute-driven disentangled representation learning, contribute to the overall performance of AD-DRL on different datasets?
4.1. Experimental Setup
In the following, we first provide the basic information of the evaluated benchmark datasets. We then detail the standard evaluation metric, followed by the key baselines used in this work, and we end this section with some implementation details.
@@ 4.1.1. Datasets
We use the widely adopted real-world recommendation dataset, the Amazon review dataset [^1] http://jmcauley.ucsd.edu/data/amazon. (McAuley and Leskovec, 2013), for evaluation in our experiments. Apart from user-item interaction data, this dataset also includes multimodal information (i.e., reviews and images) and various attributes (i.e., price, brand, etc.) of the items on 24 product categories. Three product categories from this dataset are used in the evaluation: Baby, Toys Games and Sports. Following (Liu et al., 2022), for all datasets, unpopular items and inactive users are filtered out to ensure that all the items and users have at least 5 interaction records. The basic statistics of the three datasets are shown in Table 1.
In our setting, in addition to the user-item interaction data and multimodal information of items, multiple attributes and their associated attribute values of items are required to guide the disentangled representation learning process. Specifically, we adopt four typical attributes (i.e., price, popularity  [^2] Popularity is calculated by counting the number of times each item is purchased., brand and category), and the attribute values for each attribute are compiled based on the metadata provided by the Amazon dataset: for brand and category values, we directly used the values provided by Amazon; for price and popularity, following the methods used by CoHHN (Zhang et al., 2022), we discretize them into five levels according to their numerical values. Table 1 shows the specific statistical information for each attribute in our experiments.
@@ 4.1.2. Baselines
We compared our method with the state-of-the-art methods, including both the CF-based methods (i.e., NeuMF(He et al., 2017), NGCF (Wang et al., 2019) and DGCF (Wang et al., 2020)) and Multimodal CF-based methods (JRL (Zhang et al., 2017b), MMGCN (Wei et al., 2019), MAML (Liu et al., 2019), GRCN (Wei et al., 2020), BM3 (Zhou et al., 2023) and DMRL (Liu et al., 2022)). For each type of method, we also select a disentangled representation learning model (i.e., DGCF and DMRL) for comparison.
@@ 4.1.3. Evaluation Metrics and Parameter Settings
For each dataset, we randomly split the interactions from each user with an $8:2$ ratio to construct training and testing sets. We randomly selected 10% of interactions from the training set as a validation set for hyper-parameter tuning. To evaluate the performance, we focus on the top-$n$ recommendation task, aiming to return the top-$n$ items as recommendations. We adopt two widely used evaluation metrics, i.e., Recall@$n$ and NDCG@$n$, to evaluate the accuracy, where $n$ is set as 20 by default.
The Pytorch toolkit (Paszke et al., 2019) is utilized to implement our models. To ensure a fair comparison, we optimize all the methods with Adam optimizer (Kingma and Ba, 2015), where the default learning rate and batch size are set to 0.0001 and 1024, respectively. We fixed the embedding size of each factor to 32 for AD-DRL and its variants on all datasets. The Xavier initializer (Glorot and Bengio, 2010) is used to initialize the model parameters. The $\alpha$, $\beta$, $\gamma$ and $L_{2}$ regularization coefficient are searched in $\{1e^{-5},5e^{-5},1e^{-4},\cdots,5e^{+0},1e^{+1}\}$. The number of negative examples is searched in $\{2,4,8\}$. Besides, model parameters are saved every five epochs. Early stopping strategy (Wang et al., 2019) is performed, i.e., premature stopping if Recall@20 does not increase for 50 successive epochs.
4.2. Performance Comparison (RQ1)
@@
Figure 2. Visualization of entangled and disentangled vectors from different modalities. For the second row, different colors represent different attributes. Specifically, green represents price, purple represents popularity, brown represents brand, yellow represents category, and blue represents others. For the first row, since the model does not involve attributes, different colors only represent different factors for ease of distinction.
@@
  • NeuMF, NGCF, and DGCF are the methods that are solely trained using user-item interactions. Among them, NeuMF can achieve promising performance compared to conventional MF-based methods (He et al., 2017). This is attributed to the capacity of deep neural networks to model the non-linear interactions between users and items. Benefiting from the high-order information, both NGCF and DGCF accomplish state-of-the-art results. Furthermore, DGCF performs better than NGCF by characterizing diverse user intents through the disentangled representation technique. This highlights the effectiveness of disentangled representation learning, particularly in improving the robustness of user and item representations.
  • In general, the multimodal recommendation methods perform better than those only using user-item interactions, demonstrating the effectiveness of leveraging multimodal information on learning user and item representations. Although the simple neural network structure results in lower performance of JRL compared to the graph-based methods (NDCG and DGCF), incorporating rich multi-modal information enables it to outperform NeuMF. By exploiting user-item interactions to guide the representation learning in different modalities, MMGCN yields better performance over NGCF on all the datasets. MAML outperforms NGCF by modeling users’ diverse preferences by using multimodal features of items. GRCN surpasses MAML by employing modality features to discover and prune potential false-positive edges on the user-item interaction graph. DMRL performs better than the other baselines over almost all the datasets. It should be credited to its capability to capture the different contributions of features from different modalities for each disentangled factor.
  • Furthermore, our proposed AD-DRL model consistently outperforms all baseline methods over all three datasets by a large margin. We credit this to the joint effects of the following two aspects. Firstly, robust user and item representations can be derived using attributes at different levels to guide the disentangled representation learning process. Secondly, incorporating the additional attribute information into representation learning can alleviate the data sparsity problem in recommendation.
4.3. Visualization of Disentangled Representations (RQ2)

### v2

@@ 3.2.1. High-Level Attribute-driven Disentangled Representation Learning
Intra-Modality Disentanglement. After obtaining the feature vectors of each modality, we split these vectors into several chunks. Nevertheless, different attribute factors are still entangled within the chunks. In other words, a chunk may contain both price- and brand-related features. Therefore, in order to disentangle attribute factors in each modality feature, we employ an attribute classifier to encourage each chunk to predict the corresponding attribute, as shown in Figure 1a. Taking the $k$-th chunk $\bm{v}_{t}^{k}$ of item textual embedding as an example,
(3) $\begin{cases}\bm{z}^{k}&=\bm{W}_{intra,t}\bm{v}_{t}^{k}+\bm{b}_{intra,t},\\ l_{intra,t}^{k}&=-\sum_{n=1}^{K}\tilde{z}_{n}^{k}\log\frac{\exp{z_{n}^{k}}}{\sum\nolimits_{m}\exp{z_{m}^{k}}}\end{cases}$
where $\bm{W}_{intra,t}\in\mathbb{R}^{K\times\frac{d}{K}}$ and $\bm{b}_{intra,t}\in\mathbb{R}^{K}$ are the weight matrix and bias vector of the classifier, respectively. $\bm{z}^{k}$ is the predicted logits and $z_{n}^{k}\in\bm{z}^{k}$. $\tilde{z}_{n}^{k}$ denotes the ground truth (attribute label) of chunk $\bm{v}_{t}^{k}$. For all the chunks in textual embedding $\bm{v}_{t}$, we have the following loss for all attribute factors in textual features:
@@
Inter-Modality Disentanglement. In addition to separately disentangling each modality, a serious challenge in disentangling representation learning for multi-modal features is handling the inter-relationship between the factors disentangled from multiple modalities. Intuitively, the chunks of the same attribute factor from different modality features should be consistent. For example, for the brand attribute, an item should have the same brand information in both the visual and textual features. In other words, chunks that share the same attribute across different modalities should be highly similar, while chunks that represent different attributes across modalities should be dissimilar.
To further achieve robust representation in multimodal recommendations, we disentangle attribute factors in different modality features by ensuring consistency of the same attribute across modalities. Inspired by this, we design a cross-modal contrastive loss as shown in Figure 1b. Specifically, for any two modality representations, such as $\bm{v}_{t}$ and $\bm{v}_{v}$, we take two chunks of the same attribute factor (i.e., $(\bm{v}_{t}^{k},\bm{v}_{v}^{k})$) as positive pairs, and two chunks corresponding to different attributes (i.e., $(\bm{v}_{t}^{k},\bm{v}_{v}^{n\neq k})$, $(\bm{v}_{t}^{n\neq k},\bm{v}_{v}^{k})$) as negative pairs. After that, the cross-modal contrastive loss between textual and visual features is defined as follows:
(6) $\begin{cases}l^{k}_{t\rightarrow v}&=-\log\frac{\exp((\bm{v}_{t}^{k}\cdot\bm{v}_{v}^{k})/\tau)}{\sum_{n=1}^{K}\exp((\bm{v}_{t}^{k}\cdot\bm{v}_{v}^{n})/\tau)},\\ l^{k}_{v\rightarrow t}&=-\log\frac{\exp((\bm{v}_{v}^{k}\cdot\bm{v}_{t}^{k})/\tau)}{\sum_{n=1}^{K}\exp((\bm{v}_{v}^{k}\cdot\bm{v}_{t}^{n})/\tau)},\\ l_{t\leftrightarrow v}&=\sum_{k=1}^{K}l^{k}_{t\rightarrow v}+\sum_{k=1}^{K}l^{k}_{v\rightarrow t},\end{cases}$
@@
3.2.2. Low-level Attribute-driven Disentangled Representation Learning
Each attribute $k$ is associated with a list of possible attribute values $(\tilde{y}^{k}_{1},\tilde{y}^{k}_{2},\cdots,\tilde{y}^{k}_{A_{k}})$, where $A_{K}$ is the total number of possible values for that attribute. For example, the attribute value of popularity can be stratified into five distinct levels: Super Popular, Popular, Moderate, Emerging, and Unknown. The representation disentanglement could benefit from the attribute values by exploiting their relationships of the same factor. Therefore, to learn more robust representations, we encourage disentangled representations based on attributes to predict specific attribute values of the item.
It is obvious that any single modality may not contain sufficient information for a particular attribute. For example, images may more intuitively reflect attributes such as color and brand of an item, but may not directly reflect attributes such as price or popularity. Therefore, in order to comprehensively and accurately depict the attribute values of the item, we first integrate the features from all modalities together. To achieve this, we apply a multimodal attention mechanism to measure the emphasis of different modalities on different attributes. Specifically, for the $k$-th attribute, the attention weights assigned to different modalities are estimated by a two-layer neural network:
(8) $\begin{cases}\hat{\bm{a}}^{k}=\bm{W}_{a2}\tanh(\bm{W}_{a1}(\bm{v}_{i}^{k}+\bm{v}_{t}^{k}+\bm{v}_{v}^{k})+\bm{b}_{a}),\\ \bm{a}^{k}=Softmax(\hat{\bm{a}}^{k}),\end{cases}$
where $\bm{W}_{a1}\in\mathbb{R}^{3\times\frac{d}{K}}$ and $\bm{W}_{a2}\in\mathbb{R}^{3\times 3}$ are the weight matrices corresponding to the first and second layers of the neural network, respectively. $\bm{b}_{a}$ denotes the bias vector and $\tanh$ is the activation function. $Softmax$ is adopted to normalize $\hat{\bm{a}}^{k}$ to a probability distribution.
@@ 3.2.2. Low-level Attribute-driven Disentangled Representation Learning
(9) $\bm{v}_{y}^{k}=a_{i}^{k}\cdot\bm{v}_{i}^{k}+a_{t}^{k}\cdot\bm{v}_{t}^{k}+a_{v}^{k}\cdot\bm{v}_{v}^{k},$
where $a_{i}^{k}$, $a_{t}^{k}$, $a_{v}^{k}\in\bm{a}^{k}$ are the attention weights of item ID embedding, textual feature and visual feature, respectively.
Similar to the disentangling at the intra-modality disentanglement in Section 3.2.1, we aim to achieve disentangled representation learning at the low level through attribute value prediction via a classification layer (as shown in Figure 1c):
(10) $\begin{cases}\bm{y}^{k}&=\bm{W}_{k}\bm{v}_{y}^{k}+\bm{b}_{k},\\ l^{k}_{low}&=-\sum_{i=1}^{A_{K}}\tilde{y}_{i}^{k}\log\frac{\exp{y_{i}^{k}}}{\sum\nolimits_{j}\exp{y_{j}^{k}}},\\ l_{low}&=\sum_{k=1}^{K}l^{k}_{low}\end{cases}$
where $\bm{W}_{k}\in\mathbb{R}^{K\times\frac{d}{K}}$ and $\bm{b}_{k}\in\mathbb{R}^{K}$ are the weight matrices and bias vector of the classifier, respectively. We supervise the training of each attribute subspace via independent attribute classification tasks defined in the form of a cross-entropy loss.
@@ 3.3.1. Preference Prediction
So far we have discussed how to obtain attribute-driven disentangled representations $\bm{v}_{u}=(\bm{v}_{u}^{1},\bm{v}_{u}^{2},\cdots,\bm{v}_{u}^{K})$ and $\bm{v}_{y}=(\bm{v}_{y}^{1},\bm{v}_{y}^{2},\cdots,\bm{v}_{y}^{K})$ for the user $u$ and item $i$, respectively. By assigning each disentangled representation with an attribute as described above, we are able to predict users’ preferences at the attribute level. This enhances the interpretability and controllability of our model. Specifically, to estimate a user $u$’s preference for an item $i$, it is crucial to consider her preference for each attribute of the item. To achieve this, we first compute the user’s preference score for each attribute of the item, and subsequently, integrate the scores of each attribute to estimate the user’s overall preference for the item:
(11) $\begin{cases}s_{u,i,k}=\sigma(\bm{v}_{u}^{k}\cdot\bm{v}_{y}^{k}),\\ s_{u,i}=\sum_{k=1}^{K}s_{u,i,k},\end{cases}$
@@ 3.3.2. Training Protocol
Based on the predicted preference score above, we recommend a list of top-$n$ ranked items that match the target user’s preferences. The Bayesian Personalized Ranking (BPR) loss function (Rendle et al., 2009) is employed to optimize the model parameters $\bm{\Theta}$,
@@
where $\lambda$ is the coefficient controlling $L_{2}$ regularization; $\mathcal{D}$ denotes the training set; $i_{+}$ and $i_{-}$ are the observed and unobserved items in the interaction records of user $u$, respectively. Overall, the total loss of AD-DRL is formulated as,
@@
where $\alpha$, $\beta$ and $\gamma$ are the hyperparameters to control the weight of three disentanglement modules.
@@ 4.1.1. Datasets
We use the widely adopted real-world recommendation dataset, the Amazon review dataset [^1] http://jmcauley.ucsd.edu/data/amazon. (McAuley and Leskovec, 2013), for evaluation in our experiments. Apart from user-item interaction data, this dataset also includes multimodal information (i.e., reviews and images) and various attributes (i.e., price, brand, etc.) of the items on 24 product categories. Three product categories from this dataset are used in the evaluation: Baby, Toys Games and Sports. Following (Liu et al., 2023a), for all datasets, unpopular items and inactive users are filtered out to ensure that all the items and users have at least 5 interaction records. The basic statistics of the three datasets are shown in Table 1.
In our setting, in addition to the user-item interaction data and multimodal information of items, multiple attributes and their associated attribute values of items are required to guide the disentangled representation learning process. Specifically, we adopt four typical attributes (i.e., price, popularity  [^2] Popularity is calculated by counting the number of times each item is purchased., brand and category), and the attribute values for each attribute are compiled based on the metadata provided by the Amazon dataset: for brand and category values, we directly used the values provided by Amazon; for price and popularity, following the methods used by CoHHN (Zhang et al., 2022), we discretize them into five levels according to their numerical values. Table 1 shows the specific statistical information for each attribute in our experiments. It is worth noting that our method can include a wide range of attributes that objectively reflect the character of an item, beyond the four attributes mentioned.
@@ 4.1.2. Baselines
[…]
@@ 4.1.3. Evaluation Metrics and Parameter Settings
For each dataset, we randomly split the interactions from each user into training and testing sets with an $8:2$ ratio. From the training set, 10% of interactions are randomly chosen as a validation set for tuning hyperparameters. We evaluate performance on the top-$n$ recommendation task, aiming to recommend the top-$n$ items, using Recall@$n$ and NDCG@$n$ as metrics for accuracy, with $n$ defaulting to 20.
The Pytorch toolkit (Paszke et al., 2019) is utilized to implement our models. To ensure fairness, all methods are optimized using the Adam optimizer (Kingma and Ba, 2015), with a default learning rate of 0.0001 and batch size of 1024. We fixed the embedding size of each factor to 32 for AD-DRL and its variants on all datasets. The Xavier initializer (Glorot and Bengio, 2010) is used to initialize the model parameters. The $\alpha$, $\beta$, $\gamma$ and $L_{2}$ regularization coefficient are searched in $\{1e^{-3},5e^{-3},1e^{-2},\cdots,5e^{+0},1e^{+1}\}$. The number of negative examples is searched in $\{2,4,8\}$. Besides, model parameters are saved every five epochs. Early stopping strategy (Wang et al., 2019) is performed, i.e., premature stopping if Recall@20 does not increase for 50 successive epochs. Our codes and datasets are released to facilitate the replication of our experiments  [^3] https://github.com/SDLZY/AD-DRL..
4.2. Performance Comparison
@@
  • In general, the multimodal recommendation methods perform better than those only using user-item interactions, demonstrating the effectiveness of leveraging multimodal information on learning user and item representations. Although the simple neural network structure results in lower performance of JRL compared to the graph-based methods (NDCG and DGCF), incorporating rich multi-modal information enables it to outperform NeuMF. By exploiting user-item interactions to guide the representation learning in different modalities, MMGCN outperforms NGCF on all the datasets. MAML outperforms NGCF by modeling users’ diverse preferences by using multimodal features of items. GRCN surpasses MAML by employing modality features to discover and prune potential false-positive edges on the user-item interaction graph. DMRL captures the different contributions of features from diverse modalities for each disentangled factor, achieving superior results.
  • Furthermore, our proposed AD-DRL model consistently outperforms all baselines over all three datasets by a large margin. We credit this to the joint effects of the following two aspects. Firstly, robust user and item representations can be derived using attributes at different levels to guide the disentangled representation learning process. Secondly, incorporating the additional attribute information into representation learning can alleviate the data sparsity problem in the recommendation.
4.3. Visualization of Disentangled Representations
(a) Disentangled vectors (User ID)
(b) Disentangled vectors (Item ID)
(c) Disentangled vectors (Review)
(d) Disentangled vectors (Image)
Figure 2. Visualization of disentangled vectors from ID embeddings and different modalities, with distinct colors indicating different attributes: yellow for price, orange for popularity, grey for brand, dark blue for category, and light blue for others.
AD-DRL conducts disentangled representation learning for user and item representations at two levels of attribute granularity: high and low. To further confirm the effectiveness of disentangled representation learning by AD-DRL, we use t-SNE (van der Maaten and Hinton, 2008) to cluster and visualize the disentangled vectors of users and items from the Sports dataset at each granularity level.
4.3.1. High-Level Attribute-driven Disentanglement
Figure 2 displays the disentangled vectors processed by the high-level attribute-driven disentanglement module for each modality feature, where dots of the same color denote the vectors corresponding to the same attribute. It can be observed that our model effectively distinguishes representation vectors corresponding to different attributes for both users and items across various modalities, demonstrating the effectiveness of our model in disentangling at the attribute level. Such disentangled representations can better capture user preferences and item characteristics toward different attributes.
(a) Disentangled vectors (price)
(b) Disentangled vectors (popularity)
(c) Disentangled vectors (brand)
(d) Disentangled vectors (category)
Figure 3. Visualization of the disentangled vectors corresponding to different attributes, where different colors represent different attribute values.
4.3.2. Low-Level Attribute-driven Disentanglement
To verify the effectiveness of low-level (attribute value-level) disentanglement, we visualize the representations of each attribute factor (i.e., $\bm{v}_{y}^{k}$ in Equation 9) in Figure 3. The dots of different colors represent different attribute values [^4] For brand and category, since there are too many attribute values in the dataset, making it difficult to display all of them in Figure 3. Therefore, for these two attributes, we only selected the top 5 attribute values with the most corresponding items in the dataset for demonstration.. As observed, the representations of different attribute values learned from AD-DRL are well separated and the representations of the same attribute value are concentrated. These results demonstrate that AD-DRL can achieve finer-grained disentanglement at the attribute value level, allowing AD-DRL to capture better user preferences.
4.4. Interpretability Study